# Automatically Evaluating the Paper Reviewing Capability of Large Language Models

### v2

@@ 5 Evaluation
5.1 Large Language Models
We consider eight off-the-shelf LLMs, differing in size and availability (open-source vs. proprietary): four GPT models (gpt-4o-mini, gpt-4o, o1-mini, o3-mini, o1)777gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o1-mini-2024-09-12, o1-2024-12-17, two Llama models (Llama 3.1-{70B, 405B}), and two DeepSeek models (DeepSeek-{V3, R1}). We used the default parameters of the models.
5.2 Procedure
For each of the 685 papers, we generated review data using the prompting pipeline (Section 3), extracted the strengths and weaknesses, and annotated the corresponding targets and aspects using the automatic annotator powered by o3-mini. Then for each LLM in Section 5.1, we generated a review for each paper (See Appendix A.1.2 for the prompt), extracted the strengths and weaknesses, and annotated the targets and aspects using the same automatic annotator. Then we compared the annotated targets and aspects between the experts’ reviews and LLMs’ reviews.
@@ 5.3 Result
Figure 4: Normalized distributions by target/aspect and strength/weakness for LLMs and human experts (red line). Overall, both groups showed similar perspectives in reviewing papers, focusing on technical targets (i.e., Method, Experiment, and Theory) and validity. However, LLMs showed more biased perspectives that focus on the technical validity whereas human experts exhibited more balanced perspectives. However, all the LLMs lack consideration of Novelty for weaknesses compared to human experts, which is a significant limitation in reviewing papers.
@@
While human experts raised various number of points, LLMs identified a relatively consistent number of points regardless of the paper’s content. Moreover, LLMs identified a similar number of points between strengths and weaknesses, which was a different pattern from that of the human experts. Figure 3 shows the distribution of the number of strengths and weaknesses identified by human experts and LLMs. Overall, LLMs identified more points on average (7.88) than human experts (5.39). Among the LLMs, Llama models identified fewer (3.17 strengths and 3.15 weaknesses, on average) whereas o1-mini reported more strengths and weaknesses (5.03 and 5.47, respectively) than other models. By comparing target and aspect labels between human experts and LLMs, we report the following key findings.
@@
Overall, LLMs do not effectively identify key targets and aspects when reviewing papers. Table 3 shows the overall performance of LLMs, which computes the agreement of (target, aspect) labels between human experts and LLMs. The best F1 score among the LLMs was 0.37, which indicates a low level of agreement with human experts in identifying strengths and weaknesses. Since we only considered whether the categories of review items match rather than their detailed content, the result implies that the actual content of strengths and weaknesses would be significantly different between human experts and LLMs. In general, LLMs showed higher recall than precision scores, mainly due to the nature of identifying a higher number of review points than human experts. Also, LLMs consistently achieved higher F1 scores for weaknesses than strengths. Among the LLMs, deepseek-r1 achieved the best overall performance, o1-mini achieved the best recall, and Llama-405B achieved the best precision.
@@
While overall agreement is low, both groups primarily emphasized technical validity and novelty in the strengths, and focused on technical validity and clarity in the weaknesses. Figure 4 shows the normalized distribution of target and aspect labels for both experts and LLMs. For targets, both groups primarily focused on core technical elements—Method, Experiment, and Theory. However, strengths and weaknesses illustrated different patterns: both groups praised Method more than Experiment in the strengths, but criticized Experiment more than Method in the weaknesses. For aspects, both groups considered Validity as an important aspect, especially when evaluating weaknesses. Impact received more attention than Clarity in the strengths, whereas the opposite was observed in the weaknesses.
@@
LLMs consistently exhibited a more biased perspective, while human experts maintained a more balanced perspective. Although both groups shared a core perspective, LLMs tend to focus on a few specific dimensions. For instance, LLMs focused primarily on Method and Experiment, while neglecting Prior Research (e.g., whether the paper adequately addresses prior work in positioning) and Problem (e.g., whether the task needs community attention), which human experts point out (Problem in the strengths and Prior Research in the weaknesses). For aspects, LLMs mostly focused on Validity in both strengths and weaknesses. In contrast, human experts considered the aspects more evenly among Validity, Novelty, and Clarity. Notably, LLMs exhibited a significant bias for Novelty. LLMs praised Novelty in the strengths, whereas they rarely criticized it in the weaknesses. This is a significant drawback, as a paper review requires a critical examination of novelty, by comparing them against existing work.
Table 4: F1 Score for target and aspects between DeepSeek-R1, o1-mini, and Llama-405B. Due to the biased perspective of LLMs, we observed a clear gap between what LLMs mostly focus on (e.g., Method and Experiment targets and Validity aspect) and overlook (e.g., Problem target and Novelty aspect). Full results (F1 score, precision, and recall across models and target/aspect labels) are available in Appendix A.2.4.
Target F1 score
Target DeepSeek-R1 o1-mini Llama-405B Problem 0.30 0.40 / 0.20 0.28 0.35 / 0.20 0.16 0.16 / 0.16   Method 0.73 0.75 / 0.71 0.76 0.75 / 0.77 0.69 0.76 / 0.63   Theory 0.47 0.44 / 0.51 0.47 0.41 / 0.53 0.43 0.46 / 0.40   Experiment 0.68 0.51 / 0.85 0.68 0.51 / 0.86 0.66 0.52 / 0.81
Aspect F1 score
Aspect DeepSeek-R1 o1-mini Llama-405B Novelty 0.39 0.66 / 0.12 0.39 0.66 / 0.12 0.34 0.66 / 0.01   Impact 0.41 0.54 / 0.29 0.43 0.56 / 0.30 0.32 0.35 / 0.29   Validity 0.77 0.60 / 0.95 0.77 0.60 / 0.95 0.77 0.60 / 0.95   Clarity 0.27 0.17 / 0.36 0.40 0.30 / 0.50 0.28 0.16 / 0.40
@@
Due to their biased focus, the level of agreement between LLMs and human experts varied across different labels. Table 4 shows F1 scores for specific targets and aspects. For targets and aspects that LLMs focus more on — Method and Experiment targets and Validity aspect — LLMs had a much higher level of agreement with human experts compared to other targets and aspects. In the case of Experiment, the F1 score was consistently higher for weaknesses than strengths, suggesting that LLMs are more effective at identifying concerns in experiments (e.g., lack of baselines or scope of evaluation) than recognizing strong points of theories (e.g., experiments are rigorous and thorough). Similarly, for aspects other than Validity, agreement levels were notably lower. In particular, Novelty in the weaknesses, which LLMs largely overlooked, showed a significantly lower F1 score.
@@
LLMs showed similar patterns in their emphasis in reviewing papers, regardless of their size and reasoning capability. All LLMs, including both proprietary and open source models, showed similar patterns that focused primarily on technical (Method, Experiment, and Theory) validity than on Novelty for the weaknesses. This consistency indicates that the observed biases could stem from the inherent design and training methods of LLMs, revealing potential room for improvement in the reasoning capability that requires leveraging external information (e.g., identifying comparable related work and analyzing novelty of submissions).
Final acceptance decisions are not accurate. Table 5 shows the rejection rate reported by each LLM. Overall, the best achieved rejection rate was 24.9%, which indicates that recall for rejected papers is poor. gpt-4o, Llama-405B, and DeepSeek-R1 performed significantly better than other models. gpt-4o and Llama-405B showed relatively high rejection rates while their agreement with human experts on strengths and weaknesses was low (0.348 and 0.349). DeepSeek-R1 demonstrated a moderate rejection rate among the models, with a relatively higher agreement score on strengths and weaknesses (0.373).
Table 5: Rejection percentage by model. 100% is the highest score, as we considered rejected papers. All the models were highly positive about the paper acceptance, although the papers were rejected.
Model Rejection (%)
gpt-4o 27.92%
Llama-405B 24.30%
DeepSeek-r1 23.79%
gpt-4o-mini 9.50%
DeepSeek-v3 7.93%
o1-mini 5.45%
o1 3.36%
Llama-70B 0.74%
@@ 6 Discussion
In this paper, we found gaps in the way of reviewing papers between human experts and LLMs and reported several limitations of LLMs as an automated reviewer, using an automated pipeline. Based on the results, we discuss the following implications.
@@
There exists significant room for improving alignments between human experts and LLMs in reviewing papers. Our results show that LLMs exhibit a more biased perspective, which mostly examines technical validity without contextual consideration, compared to human experts. To reduce the gap, fine-tuning models using our dataset could serve as a starting point. While our results revealed significant limitations of LLMs in reviewing papers, our focus was mostly on the target and aspect labels rather than comparing actual content. We suspect that a more significant gap lies in the actual content addressed in the review items, even if they share the same target and aspect labels. For instance, (Experiment, Validity) could point out either lack of necessary baselines or lack of ablation studies to justify authors’ arguments. Content-level investigations may reveal more limitations of LLMs, ultimately contributing to improving the reasoning capability of LLMs.
@@
Research should investigate the task of assessing the novelty of academic papers. Our finding illustrated that all LLMs in our analysis significantly overlooked the novelty aspect when evaluating weaknesses of papers. Previous studies have indicated that language models’ ability to assess novelty is inferior to that of experts Julian Just and Hutter (2024); Lin et al. (2024), emphasizing the need to encourage LLMs to focus on novelty evaluation. Although novelty is one of the most important aspects in reviewing papers and efforts have been made to enhance LLMs’ ability to assess novelty (Bougie and Watanabe, 2024; Lin et al., 2024), there exists no suitable benchmark for systematically measuring the novelty assessment capability of LLMs. We believe that creating the benchmark is a valuable contribution to the field, which allows LLMs to learn how to assess similarities between papers. Leveraging data in OpenReview could be an initial step as it contains experts’ judgement on novelty of the paper for both positive and negative decisions.
It is important to further explore the alignment between points of strengths and weaknesses and the final decision based on them. We found that there exists a discrepancy between achieving a high level of agreement in strengths and weaknesses and correctly predicting the final decision. It implies that the way LLMs make the final decision based on the identified strengths and weaknesses could be different from the way of human experts. Consistent with our findings, previous research revealed that LLMs offer positive assessments Latona et al. (2024a). As it is important to inform clear and convincing rationale behind the final decision, further investigation is needed to carefully evaluate whether the final decision based on the strengths and weaknesses is reasonable, for various stakeholders such as domain experts or novices. Generating reviews with clear alignments between the review points and final decision is a challenging task because often the relationship is not very clear and implicit Zhou et al. (2024), relying on community norms and social factors. Learning the relationship from the review data could be useful for understanding the gap between human experts and LLMs in decision-making.
@@ 7 Related Work
With the powerful reasoning capability of LLMs, LLMs have the potential to assist in the task of reviewing papers (Latona et al., 2024b; D’Arcy et al., 2024). Research has explored the capability of LLMs in reviewing papers, identifying a set of limitations. While LLM-generated reviews can be helpful (Liang et al., 2024; Tyser et al., 2024; Lu et al., 2024), research has shown that LLMs-generated reviews lack diversity (Du et al., 2024; Liang et al., 2024) and technical details (Zhou et al., 2024), exhibit bias (Ye et al., 2024), tend to provide positive feedback (Zhou et al., 2024; Du et al., 2024), and may include irrelevant or even inaccurate comments (Mostafapour et al., 2024). Furthermore, research also has reported that LLM-generated reviews have a low level of agreement with experts-generated reviews (Saad et al., 2024).
@@ 8 Conclusion
We introduced an automatic evaluation pipeline to assess LLMs’ capability in reviewing papers, by examining the agreement between LLM-generated and expert-generated reviews based on their target and aspect annotations for strengths and weaknesses. Our findings suggest that LLMs need to adopt a more balanced perspective, place greater emphasis on novelty assessment when critiquing papers, and better formulate their final judgement based on the identified strengths and weaknesses. We believe that our automated pipeline can contribute to ongoing evaluation of LLMs’ paper review capabilities within the rapid pace of LLM developments, offering concrete insights for improving their reasoning capability.
@@ Acknowledgments
This paper presents potential risks. First, while our vision is to build LLMs to effectively assist review process, our work could inadvertently encourage over-reliance on LLM-generated reviews among various user groups, including reviewers and novice researchers. Second, although our dataset could contribute to improving LLM performance of reviewing papers, it may introduce a certain bias due to the source of dataset; ICLR for papers and code based on AI research. Finally, we assess the quality of review based on alignment with expert reviews, but it could offer a potentially biased perspective, as our coding schema only considers two dimensions, which may undervalue the unique contributions of LLM-generated reviews.
@@ A.2.1 AI paper writing guidelines
A.2.2 Prompts
@@ A.2.3 Annotation Comparison
A.2.4 Results

### v3

@@ 5.1 Setup
For accepted papers, we manually collected the submitted versions of a small sample (40 papers), which has the timestamp near the ICLR deadline in the version history in arXiv. See Appendix A.5 for the focus distribution results.
@@
Models.
We consider eight off-the-shelf LLMs, differing in size and availability (open-source vs. proprietary): four GPT models (gpt-4o-mini, gpt-4o, o1-mini, o3-mini, o1)555gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o1-mini-2024-09-12, o1-2024-12-17, two Llama models (Llama 3.1-{70B, 405B}), and two DeepSeek models (DeepSeek-{V3, R1}). We also evaluated MARG D’Arcy et al. (2024) and a fine-tuned gpt-4o (see Appendix A.3 for the detail). For MARG, we only report scores for weaknesses because it only generates critiques of papers.
Metrics.
[…]
@@ 5.2 Result
Figure 4: A visualization of focus distributions by target/aspect and strength/weakness, in a descending order of cosine similarity. Overall, both groups showed similar view points in reviewing papers, focusing on technical targets (i.e., Method, Experiment, and Theory) and validity. However, LLMs showed a more biased focus towards the technical validity whereas human experts exhibited more balanced focus. Moreover, all the LLMs lack consideration of Novelty for weaknesses compared to human experts, which is a significant limitation in reviewing papers.
@@
While human experts raised various number of points, LLMs identified a relatively consistent number of points regardless of the paper’s content. Moreover, LLMs identified a similar number of points between strengths and weaknesses, which was a different pattern from that of the human experts (Figure 3). Overall, LLMs identified more points on average (7.88) than human experts (5.39). Among the LLMs, Llama models identified fewer (3.17 strengths and 3.15 weaknesses, on average) whereas o1-mini reported more strengths and weaknesses (5.03 and 5.47, respectively) than other models. The average review length of human experts and the models were 2639.76 and 3976.25, respectively. By comparing their focus distributions, we report the following key findings.
The fine-tuned gpt-4o produced focus distributions most closely aligned with that of human experts, while other models excelled in different evaluation dimensions. Table 3 shows the overall performance of the models. gpt-4o (FT) showed the highest focus distribution similarity, DeepSeek-R1 achieved the best agreement on (target, aspect) labels, and Llama-405B showed the highest text similarity score. gpt-4o showed balanced performance, with moderate scores for both focus and text similarity. The results indicate the multifaceted nature of the paper review evaluation task. In other words, assessing the quality of reviews needs a holistic approach that integrates multiple and complementary metrics.
@@
Overall, LLMs do not effectively identify key targets and aspects when reviewing papers. Table 3 shows the overall focus similarity and text similarity. The highest overall F1 score among the LLMs was 0.373, which indicates a low level of agreement with human experts in identifying strengths and weaknesses. Since we only considered whether the categories of review items match rather than their detailed content, the result implies that the actual content of strengths and weaknesses is significantly different between human experts and LLMs. In general, LLMs showed higher recall (0.402) than precision (0.300) scores, mainly due to the nature of identifying a higher number of review points than human experts. Also, LLMs consistently achieved higher F1 scores for weaknesses than strengths.
@@
While overall agreement is low, both groups have similar primary focus in reviewing papers. Figure 4 shows a visualization of focus distributions between LLMs and human experts. For targets, both groups primarily focused on core technical elements—Method, Experiment, and Theory. However, strengths and weaknesses illustrated different patterns: both groups praised Method more than Experiment in the strengths, but criticized Experiment more than Method in the weaknesses. For aspects, both groups considered Validity as the primary focus when identifying weaknesses. However, human experts focused more on Novelty in strengths whereas LLMs maintained Validity as the primary focus. For both groups, Impact received more attention in the strengths than weaknesses, whereas Clarity showed the opposite.
@@
LLMs consistently exhibited a more biased focus, notably overlooking novelty assessment in identifying weaknesses. Although both groups had the similar primary focus, LLMs tend to concentrate on a few specific dimensions. For instance, for targets, LLMs focused primarily on Method and Experiment, with less focus on Prior Research (e.g., whether the paper adequately addresses prior work in positioning) and Problem (e.g., whether the task needs community attention) compared to human experts (Problem in the strengths and Prior Research in the weaknesses). For aspects, LLMs mostly focused on Validity in both strengths and weaknesses. In contrast, human experts considered the aspects more evenly. The LLMs’ biased focus was observed for accepted papers too, mostly criticizing experimental validity (See Appendix A.5). Notably, LLMs rarely focused on Novelty aspect in identifying weaknesses. This is a significant drawback, as a paper review requires a critical examination of novelty, by comparing them against existing work. Fortunately, we observed that gpt-4o (FT) identifies Novelty aspect in the weakness, as close as human experts.
@@
Due to their biased focus, the level of agreement between LLMs and human experts varied across different labels. For targets and aspects that LLMs primarily focus on — Method (0.731, an average F1 score) and Experiment (0.671) targets and Validity (0.771) aspect — LLMs had a much higher level of agreement with human experts compared to other targets (0.213) and aspects (0.340). In the case of Experiment, the F1 score was consistently higher for weaknesses (0.835) than strengths (0.513), suggesting that LLMs are more effective at identifying concerns (e.g., lack of baselines or scope of evaluation) than strong points of experiments (e.g., experiments are rigorous and thorough). Similarly, for aspects other than Validity, agreement levels were notably lower. In particular, Novelty in the weaknesses, which LLMs largely overlooked, showed a significantly lower F1 score (0.126). See Appendix A.4 for the full results.
@@
LLMs showed similar patterns in their focus, regardless of their size and reasoning capability. All LLMs, including both proprietary and open source models, showed similar patterns that focused primarily on technical (Method, Experiment, and Theory) validity than on Novelty for the weaknesses. This consistency indicates that the observed biases could stem from the inherent design and training methods of LLMs, revealing potential room for improvement in the reasoning capability that requires leveraging external information (e.g., identifying comparable related work and analyzing novelty of submissions).
@@ 6 Discussion
In this paper, we found gaps between human experts and LLMs about their focus in reviewing papers and reported several limitations of LLMs as an automated reviewer. Based on the results, we discuss the following implications.
@@
There is significant room for improving alignments between human experts and LLMs in paper reviewing. Our results show that LLMs exhibit a more biased focus, primarily assessing technical validity without contextual consideration, compared to human experts. While fine-tuning yielded closer focus with human experts, the alignment of review points remained low. Since our focus-level evaluation only considered the target and aspect labels rather than their actual contents, we suspect that a more significant gap lies in the actual content addressed in the review items. For instance, even if two review points share the same label set (Experiment, Validity), they could point out different points such as lack of necessary baselines or lack of ablation studies to justify authors’ arguments. Content-level investigations based on annotated facets may reveal more specific limitations of LLMs in reviewing papers, ultimately contributing to improving their reasoning capability.
@@
Research should investigate the task of assessing the novelty of academic papers. Our finding illustrated that all untuned LLMs in our analysis significantly overlooked the novelty aspect when evaluating weaknesses of papers. Previous studies have indicated that language models’ ability to assess novelty is inferior to that of experts Julian Just and Hutter (2024); Lin et al. (2024), emphasizing the need to encourage LLMs to focus on novelty evaluation. Although novelty is one of the most important aspects in reviewing papers and efforts have been made to enhance LLMs’ ability to assess novelty (Bougie and Watanabe, 2024; Lin et al., 2024), there exists no suitable benchmark for systematically measuring their novelty assessment capability. We believe that creating the benchmark is a valuable contribution to the field, allowing LLMs to learn how to assess similarities between papers. Leveraging data in OpenReview could be an initial step as it contains experts’ judgment on novelty of the paper for both positive and negative decisions.
A focus-level evaluation framework can offer unique value for guiding LLM training. The automated focus-level evaluation pipeline enables continuously tracking and evaluation of how LLMs focus on key facets of a paper over time, which aligns with the goals of holistic evaluation benchmarks (Liang et al., 2022; Srivastava et al., 2022). Beyond the language model evaluation, focus-level supervision can be incorporated during the training process; reward functions can be designed to encourage balanced focus aligned with human experts or even purposefully facilitate a certain focus (e.g., building a novelty-focused reviewer) (Yang et al., 2024; Agnihotri et al., 2025). Furthermore, the framework is generalizable to other domains where the output spans multiple facets—such as debating, decision making, and educational feedback—making focus a critical factor in generated outputs.
@@ 7 Related Work
With the powerful reasoning capability of LLMs, LLMs have the potential to assist in the task of reviewing papers (Latona et al., 2024; D’Arcy et al., 2024). Research has explored the capability of LLMs in reviewing papers, identifying a set of limitations. While LLM-generated reviews can be helpful (Liang et al., 2024; Tyser et al., 2024; Lu et al., 2024), research has shown that LLMs-generated reviews lack diversity (Du et al., 2024; Liang et al., 2024) and technical details (Zhou et al., 2024), exhibit bias (Ye et al., 2024), tend to provide positive feedback (Zhou et al., 2024; Du et al., 2024), and may include irrelevant or even inaccurate comments (Mostafapour et al., 2024). Furthermore, research also has reported that LLM-generated reviews have a low level of agreement with experts-generated reviews (Saad et al., 2024).
@@ 8 Conclusion
We introduced a framework for focus-level evaluation of LLM reviews, which systematically analyzes where LLM reviews direct their praise and criticism based on pre-defined facets. Our findings suggest that LLMs need to adopt a more balanced perspective, have higher agreement with human experts about the target and aspect in the strengths and weaknesses, and place greater emphasis on novelty assessment when criticizing papers. We believe that the focus-level evaluation can contribute to ongoing evaluation of LLMs’ paper review capabilities within the rapid pace of LLM developments.
@@ Acknowledgments
This paper presents potential risks. First, while our vision is to build LLMs to effectively assist review process, our work could inadvertently encourage over-reliance on LLM-generated reviews among various user groups, including reviewers and novice researchers. Second, although our dataset could contribute to improving LLM performance of reviewing papers, it may introduce a certain bias due to the source of dataset; ICLR for papers and code based on AI research. Finally, we assess the quality of review based on alignment with expert reviews, but it could offer a potentially biased perspective, as our facets only considers two dimensions, which may undervalue the unique contributions of LLM-generated reviews.
@@ A.2.1 AI paper writing guidelines
A.2.2 Target and aspect facets
Table 5: We aim to analyze focus distributions of LLM reviews based on the targets and aspects. To identify the specific facets for targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated), we surveyed 9 AI paper submission guidelines (Appendix  A.2.1) and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). The facets were used as the codebook for human annotations.
Target
Facet Definition (The review addresses …)
Problem Motivation, task definitions, and problem statements.
Method Proposed approach, techniques, algorithms, or datasets.
Experiment Experimental setup, results, and analysis.
Conclusion Findings, implications, discussions, and takeaways.
Aspect
Facet Definition (The review addresses …)
Impact Significance or practical influence of the work.
Novelty Originality of the submission compared to prior research.
Clarity Readability, ambiguity, or communication aspects.
Validity Soundness, completeness, and rigor.
A.2.3 Prompts
@@ A.2.4 Annotation Comparison
A.3 Fine-Tuning Details
A.3.1 Fine-Tuning Dataset Construction
We constructed the fine-tuning dataset based on the corpus of papers described in Section 3. We retained 582 training samples and 98 test samples. 5 samples were excluded during tokenization due to exceeding the model’s maximum token length
A.3.2 Fine-Tuning Method
We employed supervised fine-tuning (SFT) to adapt the GPT-4o base model to our task-specific objectives. Fine-tuning was conducted using the OpenAI Fine-Tuning API [^6] https://platform.openai.com/docs/api-reference/fine-tuning, which abstracts away hardware and infrastructure details. Therefore, we do not report GPU type or compute hours. Table 6 summarizes the hyperparameter configuration used during training.
Table 6: Hyperparameter settings for supervised fine-tuning.
Parameter Value
total epochs 4
batch size 4
learning rate multiplier 0.1
A.4 Detailed Evaluation Results
@@
A.5 Results using accepted papers
Figure 17: A visualization of focus distributions by target/aspect and strength/weakness for LLMs and human experts using accepted papers, in a descending order of KL divergence. We observed a few notable differences in the pattern, compared to the evaluation results using rejected papers. First, there exists a much larger gap in the Weakness-Experiment, meaning that human experts criticize experiments significantly less than LLMs. In strengths, human experts mostly praise Novelty and Impact than Validity, but LLMs tend to praise the Validity the most. We observed the same pattern in Weakness-Novelty, meaning that LLMs neglect the novelty aspect in criticizing the papers.